A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is 20% salt and Solution B is
45% salt. She wants to obtain 150 ounces of a mixture that is 35% salt. How many ounces of each solution should she use?

Answer by Alan3354(69443)

A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is 20% salt and Solution B is
45% salt. She wants to obtain 150 ounces of a mixture that is 35% salt. How many ounces of each solution should she use?
------------
t = amount of 20%
f = amount of 45%
----
t + f = 150 (total solution)
20t + 45f = 35*150 (total salt)
-----------------
Solve for t & f